C has twice as many books as B.
If C gives B 5 books then C has 10 more books than B
How many books does C have.
I've tries this by saying that C=2b and that c-5=b+10.
I've come up with 30 but know when i solve it using 30 that's wrong.
I can figure out that the correct answer is 40, but don't know what I'm doing that I can't figure out the correct equation.

Found 2 solutions by richwmiller, LinnW:
Answer by richwmiller(17219)

c=2b,
c-5=b+5+10
2b-5=b+5+10
b=20
c=40
check
40-5=20+5+10
35=35
ok



Answer by LinnW(1048)

c=2b and that c-5=b+10
work on the second equation
add 5 to each side
c = b + 15
since c=2b and c = b + 15
2b = b + 15
add -b to each side
b = 15
c = 2*15 = 30
Test c-5 = b+10
30-5=15+10
25=25